# The Algebra of the Pseudo-Observables I: Why Quantum Mechanics is the ultimate description of Reality

### v4

@@ 2.5 Projectors
The projector 1I1-I will be called complementary of $I$. It is worth noting that if $I$ is not constant, its complementary too isn’t, besides, by virtue of (3), the product of $I$ and its complementary is null even if no one of the factors is. This demonstrates that, in the ring of observables, in general the zero-product property is not valid.
@@ 4.2 Dyad bases
where the Kronecker symbol $\delta_{l,l^{\prime}}$ is, as usual, equal to $1$ if the subscript indexes are equal and 0 otherwise. In order to do this, called $C_{jk}$ the core of the dyad $\Gamma_{jk}$, one can observe that the first condition is immediately satisfied by putting:
$C_{jj}=I_{j}$
whereas for the second, observed that, by transposition, it results:
@@
one can put:
$C_{kj}=C_{jk}^{\dagger}$
@@
where $a_{j}$ is a suitable real positive constant. We now put:
$C_{jk_{0}}=a_{j}^{-1}A_{j}\text{ .}$
So one has:
$\Gamma_{jk_{0}}\Gamma_{k_{0}j}=\Gamma_{jk_{0}}\Gamma_{jk_{0}}^{\dagger}=a_{j}^{-2}\Phi_{jk_{0}}\Phi_{jk_{0}}^{\dagger}=I_{j}=\Gamma_{jj}\text{ .}$ (44)
Observe, now, that, by virtue of (36), one has:
@@ 4.3 Change of dyad basis
We will now show as to do a change of dyad basis, i.e. as to express the dyads of a given basis $\left\{\Gamma_{jk}\right\}$, associated to the basis $\left\{I_{j}\right\}$ of pairwise orthogonal primitive projectors, in terms of the dyads of another basis $\left\{\tilde{\Gamma}_{j^{\prime}k^{\prime}}\right\}$, associated to the basis $\left\{\tilde{I}_{j^{\prime}}\right\}$ of pairwise orthogonal primitive projectors. At this aim, one has to start to choose an index $k_{0}$. Since it results:
$I_{k_{0}}=\sum_{k^{\prime}}I_{k_{0}}\tilde{I}_{k^{\prime}}\neq 0$
one may choose an index $k^{\prime}_{0}$ such that it is $I_{k_{0}}\tilde{I}_{k^{\prime}_{0}}\neq 0$. The pseudo-observable $I_{k_{0}}\tilde{I}_{k^{\prime}_{0}}$, besides, may be expressed, according the (49), as a linear combination of the dyads of the basis $\left\{\Gamma_{jk}\right\}$. By virtue of the (51) one finds that it results:
$I_{k_{0}}\tilde{I}_{k^{\prime}_{0}}=\sum_{l}\alpha_{l}\,\Gamma_{k_{0}l}$
@@
$\Gamma_{jk_{o}}\tilde{I}_{k^{\prime}_{0}}=\Gamma_{jk_{o}}\left(I_{k_{0}}\widetilde{I}_{k^{\prime}_{0}}\right)=\sum_{l}\alpha_{l}\,\Gamma_{jl}\text{ .}$
@@
$\left(\Gamma_{jk_{o}}\tilde{I}_{k^{\prime}_{0}}\right)\left(\Gamma_{kk_{o}}\tilde{I}_{k^{\prime}_{0}}\right)^{\dagger}=\sum_{l,l^{\prime}}\alpha_{l}\alpha_{l^{\prime}}^{*}\,\Gamma_{jl}\Gamma_{l^{\prime}k}=\sum_{l}\left|\alpha_{l}\right|^{2}\Gamma_{jk}$
@@
$\Gamma_{jk}=g^{2}\left(\Gamma_{jk_{o}}\tilde{I}_{k^{\prime}_{0}}\right)\left(\Gamma_{kk_{o}}\tilde{I}_{k^{\prime}_{0}}\right)^{\dagger}\text{ .}$ (61)
Decompose, now, the products $\Gamma_{jk_{o}}\tilde{I}_{k^{\prime}_{0}}$ according the dyads of the basis $\left\{\tilde{\Gamma}_{j^{\prime}k^{\prime}}\right\}$:
$\Gamma_{jk_{o}}\tilde{I}_{k^{\prime}_{0}}=\sum_{l}\beta_{lj}\,\tilde{\Gamma}_{lk^{\prime}_{0}}$ (62)
@@
$\Gamma_{jk}=\sum_{l,l^{\prime}}\left(g\,\beta_{lj}\right)\left(g\,\beta_{l^{\prime}k}\right)^{*}\tilde{\Gamma}_{lk^{\prime}_{0}}\tilde{\Gamma}_{k^{\prime}_{0}l^{\prime}}=\sum_{l,l^{\prime}}\omega_{lj}\omega_{l^{\prime}k}^{*}\,\tilde{\Gamma}_{ll^{\prime}}$ (63)
having set: $\omega_{lj}:=g\,\beta_{lj}$. Equation (63), expressing the relation among the dyads of a basis and those of another basis, may be put in a form of particular interest by introducing the pseudo-observable $\Omega$ of the change from the basis $\left\{\Gamma_{jk}\right\}$ to the basis $\left\{\tilde{\Gamma}_{jk}\right\}$:
$\Omega:=\sum_{l,m}\omega_{lm}\,\tilde{\Gamma}_{lm}\text{ .}$ (64)
@@
$\Gamma_{jk}=\Omega\tilde{\Gamma}_{jk}\Omega^{\dagger}\text{ .}$ (65)
@@
$I_{j}=\Gamma_{jj}=\Omega\tilde{\Gamma}_{jj}\Omega^{\dagger}=\Omega\tilde{I}_{j}\Omega^{\dagger}$ (66)
@@
$\tilde{\Omega}:=\sum_{j}e^{i\vartheta_{j}}\,I_{j}$ (70)
where the $\vartheta_{j}$ are real constants. It is immediate to verify that $\tilde{\Omega}$ is unitary and that the relative change of basis corresponds to one that makes it to pass from a dyad basis to an equivalent one, according to what seen at the end of subsection 4.2.

### v5

@@ 2.5 Projectors
The projector 1I1-I is the complementary projector to $I$. It is worth noting that if $I$ is not constant, its complementary too isn’t, besides, by virtue of (3), the product of $I$ and its complementary is null even if no one of the factors is. This demonstrates that, in the ring of observables, in general the zero-product property is not valid.
@@ 4.2 Dyad bases
where the Kronecker symbol $\delta_{l,l^{\prime}}$ is, as usual, equal to $1$ if the subscript indexes are equal and 0 otherwise. In order to do this, called $C_{jk}$ the core of the dyad $\Gamma_{jk}$, one can observe that, by transposition, it results:
@@
so that, for the second condition, one can put:
$C_{kj}=C_{jk}^{\dagger}\text{ .}$
@@
where $a_{j}$ is a suitable real positive constant. If we now put:
$C_{jk_{0}}=a_{j}^{-1}A_{j}$
One has:
$\Gamma_{jk_{0}}\Gamma_{k_{0}j}=\Gamma_{jk_{0}}\Gamma_{jk_{0}}^{\dagger}=a_{j}^{-2}\Phi_{jk_{0}}\Phi_{jk_{0}}^{\dagger}=I_{j}=\Gamma_{jj}$ (44)
and, therefore, the first condition is satisfied for the index $j$. Observe, now, that, by virtue of (36), one has:
@@ 4.3 Change of dyad basis
We will now show as to do a change of dyad basis, i.e. as to express the dyads of a given basis $\left\{\Gamma_{jk}\right\}$, associated to the basis $\left\{I_{j}\right\}$ of pairwise orthogonal primitive projectors, in terms of the dyads of another basis $\left\{\Gamma^{\prime}_{j^{\prime}k^{\prime}}\right\}$, associated to the basis $\left\{I^{\prime}_{j^{\prime}}\right\}$ of pairwise orthogonal primitive projectors. At this aim, one has to start to choose an index $k_{0}$. Since it results:
$I_{k_{0}}=\sum_{k^{\prime}}I_{k_{0}}I^{\prime}_{k^{\prime}}\neq 0$
one may choose an index $k^{\prime}_{0}$ such that it is $I_{k_{0}}I^{\prime}_{k^{\prime}_{0}}\neq 0$. The pseudo-observable $I_{k_{0}}I^{\prime}_{k^{\prime}_{0}}$, besides, may be expressed, according the (49), as a linear combination of the dyads of the basis $\left\{\Gamma_{jk}\right\}$. By virtue of the (51) one finds that it results:
$I_{k_{0}}I^{\prime}_{k^{\prime}_{0}}=\sum_{l}\alpha_{l}\,\Gamma_{k_{0}l}$
@@
$\Gamma_{jk_{o}}I^{\prime}_{k^{\prime}_{0}}=\Gamma_{jk_{o}}\left(I_{k_{0}}I^{\prime}_{k^{\prime}_{0}}\right)=\sum_{l}\alpha_{l}\,\Gamma_{jl}\text{ .}$
@@
$\left(\Gamma_{jk_{o}}I^{\prime}_{k^{\prime}_{0}}\right)\left(\Gamma_{kk_{o}}I^{\prime}_{k^{\prime}_{0}}\right)^{\dagger}=\sum_{l,l^{\prime}}\alpha_{l}\alpha_{l^{\prime}}^{*}\,\Gamma_{jl}\Gamma_{l^{\prime}k}=\sum_{l}\left|\alpha_{l}\right|^{2}\Gamma_{jk}$
@@
$\Gamma_{jk}=g^{2}\left(\Gamma_{jk_{o}}I^{\prime}_{k^{\prime}_{0}}\right)\left(\Gamma_{kk_{o}}I^{\prime}_{k^{\prime}_{0}}\right)^{\dagger}\text{ .}$ (61)
Let we, now, decompose the products $\Gamma_{jk_{o}}I^{\prime}_{k^{\prime}_{0}}$ according the dyads of the basis $\left\{\Gamma^{\prime}_{j^{\prime}k^{\prime}}\right\}$:
$\Gamma_{jk_{o}}I^{\prime}_{k^{\prime}_{0}}=\sum_{l}\beta_{lj}\,\Gamma^{\prime}_{lk^{\prime}_{0}}$ (62)
@@
$\Gamma_{jk}=\sum_{l,l^{\prime}}\left(g\,\beta_{lj}\right)\left(g\,\beta_{l^{\prime}k}\right)^{*}\Gamma^{\prime}_{lk^{\prime}_{0}}\Gamma^{\prime}_{k^{\prime}_{0}l^{\prime}}=\sum_{l,l^{\prime}}\omega_{lj}\omega_{l^{\prime}k}^{*}\,\Gamma^{\prime}_{ll^{\prime}}$ (63)
having set: $\omega_{lj}:=g\,\beta_{lj}$. Equation (63), expressing the relation among the dyads of a basis and those of another basis, may be put in a form of particular interest by introducing the pseudo-observable $\Omega$ of the change from the basis $\left\{\Gamma_{jk}\right\}$ to the basis $\left\{\Gamma^{\prime}_{jk}\right\}$:
$\Omega:=\sum_{l,m}\omega_{lm}\,\Gamma^{\prime}_{lm}\text{ .}$ (64)
@@
$\Gamma_{jk}=\Omega\Gamma^{\prime}_{jk}\Omega^{\dagger}\text{ .}$ (65)
@@
$I_{j}=\Gamma_{jj}=\Omega\Gamma^{\prime}_{jj}\Omega^{\dagger}=\Omega I^{\prime}_{j}\Omega^{\dagger}$ (66)
@@
$\Omega_{\vartheta}:=\sum_{j}e^{i\vartheta_{j}}\,I_{j}$ (70)
where the $\vartheta_{j}$ are real constants. It is immediate to verify that $\Omega_{\vartheta}$ is unitary and that the relative change of basis corresponds to one that makes it to pass from a dyad basis to an equivalent one, according to what seen at the end of subsection 4.2.